# Temporal Analysis of Drifting Hashtags in Textual Data Streams: A Graph-Based Application

### v1

@@ 3 Methodology
In this section, the data collection is detailed, discussed, and the obtained dataset is described. In addition, we present the hashtags treatment.
@@ 3.1 Dataset description
The third peak, between May and August 2022, reflects an important event regarding the #mybodymychoice. In June 2022, the US Supreme Court invalidated the Roe vs. Wade decision (which favored abortion in 1973) [^8] Available at: https://www.supremecourt.gov/opinions/21pdf/19-1392_6j37.pdf. Accessed on May 30th, 2023..
In summary, the data collected provides several insights and reflects the past regarding the hashtag #mybodymychoice. The link for downloading the dataset and source code will be made upon paper acceptance.
3.2 Hashtags treatment
@@ 4.1 Graph generation
We generated graphs using their intrinsic characteristics: each node corresponded to a hashtag, and each edge between two nodes corresponded to their co-occurrence. We considered a hashtag valid once it appears in the dataset at least five times to ease the analyses and avoid noise. Considering the textual data stream setting, we perform the computation on-the-fly. It means that the graph generation process respects the textual data stream constraints. The only exception is the community detection performed at the end of each year (regarding the timestamp of the textual data stream).
In addition, to limit our analysis to the most recent and important hashtags, we used a 200-size window and a hashtag aging mechanism. In our approach, each hashtag has an age attribute. This attribute controls the hashtags to be considered in the analysis. Whenever a new hashtag is to be inserted, if the window is full, the oldest hashtag in the window is removed to accommodate the new hashtag. If the window is not full, the hashtag is added. In each iteration, the hashtags’ age is increased, except for the recently received hashtag.
@@ 4.2 Implementation
We implemented the code in Python 3.9. We plotted the graphs with the help of the Pyvis [^9] https://pyvis.readthedocs.io/en/latest/ and NetworkX libraries [^10] https://networkx.org/. The link for downloading the source code and dataset will be available upon paper acceptance. Algorithm 1 depicts the pseudocode. The algorithm receives a textual data stream $TS$ and creates a new Graph with a window size of 200 and a minimum frequency of five. Line 2 starts the stream iteration; line 3 obtains the next text (tweet); line 4 extracts a list of hashtags, which is sent to the add_node graph’s method, detailed in Algorithm 2.
@@
The Algorithm 2 details the add_node method. In line 1, all the hashtags’ age is incremented. In line 3, for each hashtag in the received list, we check if the hashtag is not in the vocabulary (vocab), i.e., it is not in the graph, if it is in the pregraph, and if its frequency reached five simultaneously. Pregraph is a graph’s internal structure that stores the hashtags until they reach the minimum defined frequency. Also, g.vocab is the structure in which we control the existence of nodes, i.e., hashtags. Once a given hashtag reaches the minimum frequency, it can be added to the graph in line 4. In line 5, the hashtag is removed from the pregraph. If the condition in line 3 is not met, it is checked if the hashtag is in the pregraph. If the condition is met, its frequency is incremented. Next, we evaluate the remaining hashtags as connections. In line 10, we call the variable other_hashtags. The same process of checking existence in g.vocab and pregraph is performed between lines 12 and 18. In line 19, we connect the hashtag and the other_hashtag. In lines 20 and 21, we reset the age of hashtag and other_hashtag. The age reset is important so the window can identify and remove stale hashtags when opportune.
@@
It is important to mention that we do not include the hashtag #mybodymychoice because all the nodes would be connected to the #mybodymychoice’s node.
@@ 5.1 Analysis of 2018
Figure 4: Part of the green community from 2018.
Figure 5: Part of the purple community from 2018.
@@
Figure 6: Part of the orange community from 2018.
Figure 7: Part of the yellow community from 2018.
@@
In summary, regarding 2018, the hashtag #mybodymychoice was used in different contexts than its original one, drifting to civil rights, the use of drugs, and politics. However, the largest community regards the original context: feminism, abortion, and women’s bodily autonomy. The five most frequent hashtags in 2018 and respective frequencies were: #prochoice (987), #repealthe8th (532), #womensrights (482), #metoo (379), and #fbr (339).
@@ 5.2 Analysis of 2019
Figure 10: Part of the purple community from 2019.
The community in blue also stands out. It contains hashtags such as #mykidsmychoice, #votethemout, #firstamendment, #votenos2173, and #killbills2173. The hashtag #votethemout is probably related to the unacceptance of the government in charge in 2019. The hashtag #firstamendment prevents the US government from making laws prohibiting religions and their exercise [^12] Available at: https://www.whitehouse.gov/about-the-white-house/our-government/the-constitution. Accessed on May 27th, 2023.. In addition, #killbills2173 and #votenos2173 regard Bill S2173131313Available at: https://www.billtrack50.com/BillDetail/966141. Accessed on May 31st, 2023., which intended to eliminate the religious exemption for mandatory childhood vaccination. Therefore, this community combines hashtags related to religion, politics, and vaccination. Fig. 11 shows the major part of the community in blue.
Figure 11: Part of the blue community from 2019.
@@
Figure 12: Part of the blue-gray community from 2019.
Figure 13: Part of the dark-orange community from 2019.
It is also possible to notice that new communities are growing in 2019 but still have a few components. In addition, we see that the #mybodymychoice regards vaccination, feminism, civil rights, and politics in this time slice. The five most frequent hashtags in 2019 and respective frequencies were: #prochoice (2451), #womensrights (1384), #womensrightsarehumanrights (1207), #abortionisawomansright (916), and #abortion (905).
@@ 5.3 Analysis of 2020
Fig. 14 depicts the communities discovered by the Girvan-Newman algorithm considering the hashtags until the end of 2020. Two big communities draw attention: those in blue and yellow.
@@
Exploring the biggest community, i.e., the blue-colored community, we see hashtags such as #abortion, #abortionrights, #reproductivehealth, #womensrights, #abortionishealthcare, #abortionwithoutborders, and #abortionisnotacrime. In addition, hashtags in languages other than English expressing the same stance are present, i.e., #abortolegal (“legal abortion” in Portuguese or Spanish), #avortement (“abortion” in French), and #abortolegalya (“legal abortion now” in Spanish). However, since #abortion is present in this community, opinions against abortion are present, for example, the hashtags #rescuethepreborn, #chooselife, #savethebabyhumans, and #endabortions. Therefore, this community, depicted in Fig. 15, combines hashtags in favor and against abortion.
@@
In November 2019, a new pneumonia-related disease appeared in China, later named Covid-19, after coronavirus disease and the year this disease emerged. Due to its ease to spread and severity, the World Health Organization (WHO) declared Covid-19 a global pandemic in March 2020 Cucinotta and Vanelli (2020). Therefore, WHO suggested measures to prevent the Covid-19 spread, such as wearing masks, keeping distance from other people, and washing hands [^14] Available at: https://www.who.int/westernpacific/emergencies/covid-19/information/transmission-protective-measures. Accessed on May 30th, 2023.. The US Food and Drugs Association (FDA) authorized the first Covid-19 vaccine for emergency use in December 2020151515Available at: https://www.fda.gov/news-events/press-announcements/fda-takes-key-action-fight-against-covid-19-issuing-emergency-use-authorization-first-covid-19. Accessed on May 30th, 2023..
@@
In summary, although there are hashtags related to the original aim of #mybodymychoice, it is possible to see that the hashtag #mybodymychoice shifted to topics other than abortion and women’s bodily autonomy. Covid-19-related hashtags emerged in 2020. The five most frequent hashtags in 2020 and respective frequencies were: #prochoice (134), #abortion (64), #strajkkobiet (54), #votenos2173 (53), and #prolife (50).
@@ 5.4 Analysis of 2021
The hashtags relationships collected at the end of 2021 are depicted in Figure 19. The community in yellow contains hashtags such as #dictator, #eucovidcertificate, #antivax, #novaccinepassportanywhere, #fighttyrany, #freedomofchoice, #novax, and #covidscam. We see that most hashtags are against vaccination, vaccine passport, and in favor of having the choice to have the vaccine or not. Although these hashtags essentially express opinions against vaccination or its governmental control, some hashtags in favor of vaccination also appear. For example, #getvaccinated and #getvaccinatednow appear among the contrary hashtags. Considering the analysis of the studied period, i.e., between 2018 and 2022, this is the first time the biggest community is not related to abortion and women’s bodily autonomy. Fig. 20 shows part of the yellow community.
@@
Figure 20: Part of the yellow community from 2021.
@@
Figure 22: Part of the dark-orange community from 2021.
Figure 23: Part of the blue-gray community from 2021.
In summary, the analysis of 2021 highlights the first time in the analyzed period that the biggest detected community does not relate to abortion and women’s bodily autonomy. It shows that, at least temporarily, the hashtag #mybodymychoice drifted to vaccination and Covid-19-related topics. The five most frequent hashtags in 2021 and respective frequencies were: #novaccinepassports (3694), #prochoice (2736), #novaccinepassportsanywhere (2605), #freedom (2524), and #novaccinemandates (2305).
@@ 5.5 Analysis of 2022
Figure 25: Part of the blue community from 2022.
The community in yellow contains hashtags related to Covid-19. For example, #novaccinemandates, #vaccineinjuries, #vaccinesideeffects, #informedconsent, #standupforfreedom, #vaccinedisaster, #lockdowns, and #bigpharma appear in this community. Different from 2021, in 2022, hashtags related to vaccine side effects and injuries appear more frequently than hashtags against vaccination in a Covid-19-related group. The dark-green community also considers Covid-19-related hashtags such as #greenpass, #nojabnojob, #italianblackmail, #novax, and #pfizergate. It is closely related to Covid-19 vaccination since Green Pass is the name on which the EU Digital Covid Certificate is known in Italy [^17] Available at: https://italygreenpass.com/. Accessed on May 30th, 2023.. A dark-orange community also emerged containing hashtags related to Covid-19, such as #yourbodyyourchoice, #thejab, and #covidisntover. These communities are partially depicted in Figs. 26, 27, and 28 respectively.
Figure 26: Part of the yellow community from 2022.
Figure 27: Part of the dark-green community from 2022.
Figure 28: Part of the dark-orange community from 2022.
The community in blue-gray regards sexual topics, such as #nudes, which consider nudity. Also, the community in dark red is similar to the dark-orange community, having hashtags such as #braless and #freethenipples. The community in light orange combines different hashtags, e.g., #mutilation, #genital, #cutting, #children, and #circumcision. This last community is related to the circumcision topic, probably against it, since other hashtags, such as #bornperfect, #trauma, and #ididnotconsent are also present, as can be seen in Fig. 29.
@@
In summary, the biggest communities in the analysis of 2022 regard mostly Covid-19, abortion, and women’s bodily autonomy-related topics. Other topics, such as politics and sexual-related, also appear. We also highlight that hashtags related to the Russian-Ukrainian conflict and vaccination side effects emerged. The five most frequent hashtags in 2022 and respective frequencies were: #roevswade (17719), #roevwade (11179), #womensrights (8750), #prochoice (7987), and #womensrightsarehumanrights (7961).

### v2

@@ 3 Methodology
In this section, the data collection is detailed and discussed, and the obtained dataset is described. In addition, we present the treatment of hashtags.
@@ 3.1 Dataset description
The third peak, between May and August 2022, reflects a significant event regarding the #mybodymychoice. In June 2022, the US Supreme Court invalidated the Roe vs. Wade decision (which favored abortion in 1973) [^8] Available at: https://www.supremecourt.gov/opinions/21pdf/19-1392_6j37.pdf. Accessed on May 30th, 2023..
In summary, the data collected provides several insights and reflects the past regarding the hashtag #mybodymychoice. Under Twitter’s policies, the data and files with the generated graphs are publicly available at https://github.com/cristianomg10/temporal-analysis-of-drifting-hashtags-in-textual-data-streams-a-graph-based-application.
3.2 Treatment of Hashtags
@@
3.3 Evaluation
The evaluation of the approach is performed qualitatively. This means that the analysis is performed after each year finishes (according to the tweets’ timestamps). The most extended communities have their nodes, i.e., hashtags, analyzed in order to find patterns and infer the context.
Since we analyze the hashtag #mybodymychoice in this paper, we consider drifts to have occurred whenever the hashtag is used in contexts different from its original. That is, we consider drift if a community’s context is different from feminism and bodily autonomy. In addition, in case of a drift occurrence, we consider the size of the community as the drift importance.
@@ 4.1 Graph generation
We generated graphs using their intrinsic characteristics: each node corresponded to a hashtag, and each edge between two nodes corresponded to their co-occurrence. We considered a hashtag valid once it appears in the dataset at least five times to ease the analyses and avoid noise. We perform the computation on the fly considering the textual data stream setting. This means that the graph generation process respects the constraints of the textual data stream. The only exception is the community detection performed at the end of each year (regarding the timestamp of the textual data stream).
In addition, to limit our analysis to the most recent and important hashtags, we used a 200-size window and a hashtag aging mechanism. In our approach, each hashtag has an age attribute. This attribute controls the hashtags to be considered in the analysis. Whenever a new hashtag is to be inserted, if the window is full, the oldest hashtag in the window is removed to accommodate the latest hashtag. If the window is not full, the hashtag is added. In each iteration, the hashtags’ age increases, except for the recently received hashtag.
@@ 4.2 Implementation
We implemented the code in Python 3.9. We plotted the graphs with the help of the Pyvis [^9] https://pyvis.readthedocs.io/en/latest/ and NetworkX libraries [^10] https://networkx.org/. Algorithm 1 depicts the pseudocode. The algorithm receives a textual data stream $TS$ and creates a new Graph with a window size of 200 and a minimum frequency of five. Line 2 starts the stream iteration; line 3 obtains the following text (tweet); line 4 extracts a list of hashtags, which is sent to the add_node graph’s method, detailed in Algorithm 2.
@@
The Algorithm 2 details the add_node method. In line 1, all the hashtags’ age is incremented. In line 3, for each hashtag in the received list, we check if the hashtag is not in the vocabulary (vocab), i.e., it is not in the graph, if it is in the pregraph, and if its frequency reached five simultaneously. Pregraph is a graph’s internal structure that stores the hashtags until they reach the minimum defined frequency. Also, g.vocab is the structure in which we control the existence of nodes, i.e., hashtags. Once a given hashtag reaches the minimum frequency, the graph accommodates it in line 4. In line 5, the hashtag is removed from the pregraph. If the condition in line 3 is not met, it is checked if the hashtag is in the pregraph. If the condition is met, its frequency is incremented. Next, we evaluate the remaining hashtags as connections. In line 10, we call the variable other_hashtags. The same process of checking existence in g.vocab and pregraph is performed between lines 12 and 18. In line 19, we connect the hashtag and the other_hashtag. In lines 20 and 21, we reset the age of hashtag and other_hashtag. The age reset is crucial so the window can identify and remove stale hashtags when opportune.
@@
We do not include the hashtag #mybodymychoice because all the nodes would connect to the #mybodymychoice’s node, hampering the visualization and including an unnecessary overhead to the algorithm.
The computational complexity of our approach is $O(SH^{2})$, where $H$ is the hashtags of a given tweet (post), and $S$ represents the stream. We consider the $H^{2}$ component the complexity of the relationship mapping between the hashtags of a tweet. Please notice that we use the text stream paradigm. Therefore, we process tweet by tweet. In this point of view, the complexity is $O(H^{2})$.
4.3 Software and Hardware specification
The experiment and analysis described in this paper were performed using Python [^11] https://www.python.org/, version 3.9.12. In addition, we used Pyvis [^12] https://pyvis.readthedocs.io/en/latest/ for graph visualization and NetworkX [^13] https://networkx.org/ for the implementation of the Girvan-Newman method. We also used Matplotlib [^14] https://matplotlib.org/ and Seaborn [^15] https://seaborn.pydata.org/ to generate some of the figures for this paper. The experiment and analysis were run on a laptop with a 12th-generation Intel 2.3GHz and 64 GB of RAM on Windows 11.
@@ 5.1 Analysis of 2018
Figure 4: Partial view of the green community from 2018.
Figure 5: Partial view of the purple community from 2018.
@@
Figure 6: The orange community from 2018.
Figure 7: Yellow community from 2018.
@@
In summary, regarding 2018, the hashtag #mybodymychoice was used in different contexts than its original one, drifting to civil rights, the use of drugs, and politics. However, the largest community regards the original context: feminism, and women’s bodily autonomy. The five most frequent hashtags in 2018 and respective frequencies were: #prochoice (987), #repealthe8th (532), #womensrights (482), #metoo (379), and #fbr (339). Table 1 summarizes the top 5 communities detected at the end of 2018 sorted by size, with their respective hashtag counts.
Context # of Hashtags
Feminism and bodily autonomy 90
Mixed 68
Civil rights 17
Mixed 12
Drug liberation 7
Table 1: Top 5 communities and their respective contexts and number of hashtags, for 2018.
@@ 5.2 Analysis of 2019
Figure 10: Partial view of the purple community from 2019.
The community in blue also stands out. It contains hashtags such as #mykidsmychoice, #votethemout, #firstamendment, #votenos2173, and #killbills2173. The hashtag #votethemout is probably related to the unacceptance of the government in charge in 2019. The hashtag #firstamendment prevents the US government from making laws prohibiting religions and their exercise [^17] Available at: https://www.whitehouse.gov/about-the-white-house/our-government/the-constitution. Accessed on May 27th, 2023.. In addition, #killbills2173 and #votenos2173 regard Bill S2173181818Available at: https://www.billtrack50.com/BillDetail/966141. Accessed on May 31st, 2023., which intended to eliminate the religious exemption for mandatory childhood vaccination. Therefore, this community combines hashtags related to religion, politics, and vaccination. Fig. 11 shows the major part of the community in blue.
Figure 11: Partial view of the blue community from 2019.
@@
Figure 12: Partial view of the blue-gray community from 2019.
Figure 13: Partial view of the dark-orange community from 2019.
It is also possible to notice that new communities are growing in 2019, but they still have a few components. In addition, we see that the #mybodymychoice regards vaccination, feminism, civil rights, and politics in this time slice. The five most frequent hashtags in 2019 and respective frequencies were: #prochoice (2451), #womensrights (1384), #womensrightsarehumanrights (1207), #abortionisawomansright (916), and #abortion (905). Table 2 summarizes the top 5 communities detected at the end of 2019 with their respective hashtag counts.
Context # of Hashtags
Civil rights, women’s rights, and campaign topics 86
Anti-abortion 28
Religion, politics, and vaccination 26
Feminism/Hate speech 16
Vape liberation 11
Table 2: Top 5 communities and their respective contexts and number of hashtags for 2019.
@@ 5.3 Analysis of 2020
Fig. 14 depicts the communities discovered by the Girvan-Newman algorithm considering the hashtags until the end of 2020. Two significant communities draw attention: those in blue and yellow.
@@
Exploring the most extensive community, i.e., the blue-colored community, we see hashtags such as #abortion, #abortionrights, #reproductivehealth, #womensrights, #abortionishealthcare, #abortionwithoutborders, and #abortionisnotacrime. In addition, hashtags in languages other than English expressing the same stance are present, i.e., #abortolegal (“legal abortion” in Portuguese or Spanish), #avortement (“abortion” in French), and #abortolegalya (“legal abortion now” in Spanish). However, since #abortion is present in this community, opinions against abortion are present, for example, the hashtags #rescuethepreborn, #chooselife, #savethebabyhumans, and #endabortions. Therefore, this community, depicted in Fig. 15, combines hashtags in favor and against abortion.
@@
In November 2019, a new pneumonia-related disease appeared in China, later named Covid-19, after coronavirus disease and the year this disease emerged. Due to its ease of spread and severity, the World Health Organization (WHO) declared Covid-19 a global pandemic in March 2020 [Cucinotta & Vanelli, 2020]. Therefore, WHO suggested measures to prevent the Covid-19 spread, such as wearing masks, keeping distance from other people, and washing hands [^19] Available at: https://www.who.int/westernpacific/emergencies/covid-19/information/transmission-protective-measures. Accessed on May 30th, 2023.. The US Food and Drugs Association (FDA) authorized the first Covid-19 vaccine for emergency use in December 2020202020Available at: https://www.fda.gov/news-events/press-announcements/fda-takes-key-action-fight-against-covid-19-issuing-emergency-use-authorization-first-covid-19. Accessed on May 30th, 2023..
@@
In summary, although there are hashtags related to the original aim of #mybodymychoice, it is possible to see that the hashtag #mybodymychoice shifted to topics other than abortion and women’s bodily autonomy. Covid-19-related hashtags emerged in 2020. The five most frequent hashtags in 2020 and respective frequencies were: #prochoice (134), #abortion (64), #strajkkobiet (54), #votenos2173 (53), and #prolife (50). Table 3 summarizes the top 5 communities detected at the end of 2020 with their respective hashtag counts and contexts.
Context # of Hashtags
Abortion, and women’s rights 111
Vaccination, and Covid-19 vaccination 57
Against governmental measures regarding Covid-19 11
Circumcision 6
Not possible to infer 2
Table 3: Top 5 communities and their respective contexts and number of hashtags for 2020.
@@ 5.4 Analysis of 2021
The hashtags relationships collected at the end of 2021 are depicted in Figure 19. The community in yellow contains hashtags such as #dictator, #eucovidcertificate, #antivax, #novaccinepassportanywhere, #fighttyrany, #freedomofchoice, #novax, and #covidscam. We see that most hashtags are against vaccination and vaccine passport, and in favor of having the choice to have the vaccine or not. Although these hashtags essentially express opinions against vaccination or its governmental control, some hashtags in favor of vaccination also appear. For example, #getvaccinated and #getvaccinatednow appear among the contrary hashtags. Considering the analysis of the studied period, i.e., between 2018 and 2022, this is the first time the most extensive community is not related to abortion and women’s bodily autonomy. Fig. 20 shows part of the yellow community.
@@
Figure 20: Partial view of the yellow community from 2021.
@@
Figure 22: Dark-orange community from 2021.
Figure 23: Blue-gray community from 2021.
In summary, the analysis of 2021 highlights the first time in the analyzed period that the biggest detected community does not relate to abortion and women’s bodily autonomy. It shows that, at least temporarily, the hashtag #mybodymychoice drifted to vaccination and Covid-19-related topics. The five most frequent hashtags in 2021 and respective frequencies were: #novaccinepassports (3694), #prochoice (2736), #novaccinepassportsanywhere (2605), #freedom (2524), and #novaccinemandates (2305). Table 4 summarizes the top 5 communities detected at the end of 2021 with their respective hashtag counts and contexts.
Context # of Hashtags
Covid-19, vaccination, and governmental measures 123
Abortion, and women’s bodily autonomy 45
Against Covid-19 vaccines 14
Against lockdown and Trudeau (Canadian Prime Minister in 2021) 4
Not possible to infer 2
Table 4: Top 5 communities and their respective contexts and number of hashtags for 2021.
@@ 5.5 Analysis of 2022
Figure 25: Partial view of the blue community from 2022.
The community in yellow contains hashtags related to Covid-19. For example, #novaccinemandates, #vaccineinjuries, #vaccinesideeffects, #informedconsent, #standupforfreedom, #vaccinedisaster, #lockdowns, and #bigpharma appear in this community. Different from 2021, in 2022, hashtags related to vaccine side effects and injuries appear more frequently than hashtags against vaccination in a Covid-19-related group. The dark-green community also considers Covid-19-related hashtags such as #greenpass, #nojabnojob, #italianblackmail, #novax, and #pfizergate. It is closely related to Covid-19 vaccination since Green Pass is the name on which the EU Digital Covid Certificate is known in Italy [^22] Available at: https://italygreenpass.com/. Accessed on May 30th, 2023.. A dark-orange community also emerged containing hashtags related to Covid-19, such as #yourbodyyourchoice, #thejab, and #covidisntover. Figs. 26, 27, and 28 respectively depict partially these communities.
Figure 26: Partial view of the yellow community from 2022.
Figure 27: Partial view of the dark-green community from 2022.
Figure 28: Partial view of the dark-orange community from 2022.
The community in blue-gray regards sexual topics, such as #nudes, which is considered nudity. Also, the community in dark red is similar to the dark-orange community, having hashtags such as #braless and #freethenipples. The community in light orange combines different hashtags, e.g., #mutilation, #genital, #cutting, #children, and #circumcision. This last community is related to the circumcision topic, probably against it, since other hashtags, such as #bornperfect, #trauma, and #ididnotconsent are also present, as can be seen in Fig. 29.
@@
In summary, the most extensive communities in the analysis of 2022 regard mostly Covid-19, abortion, and women’s bodily autonomy-related topics. Other topics, such as politics and sexual-related, also appear. We also highlight that hashtags related to the Russian-Ukrainian conflict and vaccination side effects emerged. The five most frequent hashtags in 2022 and respective frequencies were: #roevswade (17719), #roevwade (11179), #womensrights (8750), #prochoice (7987), and #womensrightsarehumanrights (7961). Table 5 summarizes the top 5 communities analyzed at the end of 2022 with their respective hashtag counts and contexts.
Context # of Hashtags
Women’s bodily autonomy 62
Covid-19 vaccination 40
Sex-related topics 15
Nudity 14
Circumcision 9
Table 5: Top 5 communities and their respective contexts and number of hashtags for 2022.